OpenStudy (anonymous):

am i right?!? What is the sum of the first 19 terms of the series: 63 + 55 + 47 + 39 + 31 + …? a. -2565 b. -171 c. 247 d. 769.5 My answer: b. -171

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow {n \over 2}(a_1 + a_n) }\)

OpenStudy (anonymous):

@ParthKohli so was my answer correct??

OpenStudy (campbell_st):

the formula when you don't have the last term is \[s _{n} = \frac{n}{2}[2a + n-1)d]\] a = 1st term, d = common difference and n = number of terms in your question a = 63, d = -8 and n = 19 substitute and evaluate

OpenStudy (ash2326):

Yeah you're right

Parth (parthkohli):

First find a19. \(\Large \color{Black}{\Rightarrow a_{19} = a_1 + -8(19 - 1) }\) \(\Large \color{Black}{\Rightarrow 63 + -8(18) }\) \(\Large \color{Black}{\Rightarrow 63 - 144 }\) \(\Large \color{Black}{\Rightarrow -81 }\)

Parth (parthkohli):

And there you go correct with the formula

OpenStudy (campbell_st):

oops n-1 sohould read (n -1)

OpenStudy (anonymous):

Yay! so i'm right?!?!

jimthompson5910 (jim_thompson5910):

You first need the 19th term, which is a19 = a1+(n-1)*d a19 = 63+(19-1)*(-8) a19 = -81 Then use it in the formula Sn = (n*(a1+an))/2 S19 = (19*(63+(-81)))/2 S19 = -171

jimthompson5910 (jim_thompson5910):

oops, should have put the line S19 = (19*(a1+a19))/2

OpenStudy (anonymous):

awesome! thank you!!! :) can you please give @ParthKohli a medal @jim_thompson5910 cuz this only lets me give one out and i want to give both of you one! :D

jimthompson5910 (jim_thompson5910):

you're welcome